OpenStudy (anonymous):

What is the approximate circumference of a circle whose radius is 6 feet long? Round your answer to one decimal place.

OpenStudy (anonymous):

I got 18.8

OpenStudy (midhun.madhu1987):

Circumference of a Circle of radius "r" is given by: C = 2* pi * r pi = 3.14

OpenStudy (midhun.madhu1987):

C = 2 * 3.14 * 6 = ?

OpenStudy (anonymous):

37.68

OpenStudy (midhun.madhu1987):

Yes.. Perfect.. But you need to round it off to one place....

OpenStudy (anonymous):

37.7

OpenStudy (anonymous):

thank you

OpenStudy (midhun.madhu1987):

yw :)